OpenStudy (anonymous):

Find f(x)·g(x) if f(x) = 3x and g(x) = x + 1. A. 3x^2 + 3x B. 3x + 1 C. 4x + 1 D. 3x^2 + 1

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Multiply 3x and x + 1 together.

OpenStudy (anonymous):

(3x)·(x+1)

OpenStudy (anonymous):

A right?

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

Find F o G if f(x) = 2x - 1 and g(x) = x + 2. A. 2x - 3 B. 2x + 1 C. 2x - 1 D. 2x + 3

OpenStudy (anonymous):

(F o G)(x) = f( g(x) ) so if f(x) = 2x - 1 and g(x) = x + 2, then f( g(x) ) = 2(x+2) - 1 Does this make sense?

OpenStudy (anonymous):

I got B

OpenStudy (anonymous):

2(x+2) - 1 = 2x + 4 - 1 = 2x + 3

OpenStudy (anonymous):

Is this clear? No?

OpenStudy (anonymous):

kinda My answer is D.

OpenStudy (anonymous):

Yeah. What part don't you understand? I'd be glad to clarify it.

OpenStudy (anonymous):

for the second one yea it's D

OpenStudy (anonymous):

so the answer is D right?

OpenStudy (anonymous):

Yes.